plz answer all and give exolination thanks
4. The magnetic field inside a superconducting solenoid is 4.70 T. The solenoid has an inner diameter of 6.20 cm and a length of 26.0 cm.
(a) Determine the magnetic energy density in the field.
uB = J / m3
(b) Determine the energy stored in the magnetic field within the solenoid.
UB = kJ

5.–/2 pointsPSE6 32.P.032.
At t = 0, an emf of 540 V is applied to a coil that has an inductance of 0.820 H and a resistance of 30.0 .
(a) Find the energy stored in the magnetic field when the current reaches half its maximum value.
J
(b) After the emf is connected, how long does it take the current to reach this value?
ms

6.–/1 points PSE6 32.P.046.
A 0.95 µF capacitor is charged by a 44.0 V power supply. The fully-charged capacitor is then discharged through a 9.5 mH inductor. Find the maximum current in the resulting oscillations.
Answer #1

4 a) Let us first calculate the volume of the solenood which is given by = π d2/4 × L = 7.84 × 10-4 m3.

Magnetic energy is given by the formula = B2/2uo

B = u0 n I

Hence magnetic energy =uo n2 I2/2= 4.24 × 10-8 J

Hence Magnetic energy density = Magnetic Energy /Volume

= 5.4 × 10-5 J/m3

b) Energy stored in the magnetic field = 4.24 × 10-8 J

2) Maximum current value = E/R = 540/30 A = 18 A

a)Energy stored in the magnetic field = 1/2 L I^2 = 33.21 J

b)

When current is half of maximum value

Then e ​​​​​​-Rt/L = 1/2

Hence t = 0.089 sec

c) Maximum current in LC oscillation is given by

wq0 Where w =√(1/LC) and qo = maximum charge = CE

Max current = 1.04 × 10-5 ×4.04 × 10^(-5)

4.2 × 10^(-10) A